OpenStudy (anonymous):

The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

I think it's obvious ,right?

OpenStudy (anonymous):

\[slope of any line m =\frac{ 5-6 }{5-\left( -1 \right) }=\frac{ -1 }{6 }\] slope of line perpendicular to the given line=-1/m=6 \[\mid point of R and S=(\frac{ -1+5 }{2 },\frac{ 6+5 }{2 } )=( 2,\frac{ 11 }{2 } )\] Eq. of any line through (2,11/2 ) and slope 6 is \[y-\frac{ 11 }{2 }=6\left( x-2 \right)\] 2y-11=6x-12 or 2 y-6 x+1=0

OpenStudy (anonymous):

okay, I think I understand the process of calculating. thank you!

OpenStudy (anonymous):

bisector of R and S= midpoint of R and S